OpenStudy (anonymous):

use implicit differentiation to find an equation of the tangent line to the curve at the given point, xsin2y=ycos2x, (pi/2,pi/4) Thanks

OpenStudy (bahrom7893):

\[x*Cos(2y)*2\frac{dy}{dx}+Sin(2y)=y*(-Sin(2x))*2+Cos(2x)*\frac{dy}{dx}\]

OpenStudy (bahrom7893):

(using the product and chain rules)

OpenStudy (bahrom7893):

Now just plug in the values you know and solve for \[\frac{dy}{dx}\]

OpenStudy (anonymous):

so i'm plugging in pi/2 for x and pi/4 for y?

OpenStudy (bahrom7893):

yup. And then solving for \[\frac{dy}{dx}\]

OpenStudy (anonymous):

i got dy/dx=-1

OpenStudy (bahrom7893):

You're probably right. Sorry, I don't feel like plugging it all in, but if you post your work, I'll check it over for you.

OpenStudy (anonymous):

i feel confident thank you though for your help